# Lost in Overlap: Exploring Logit-based Watermark Collision in LLMs

### v2

@@ Abstract
Introduction
As the quality of text produced by large language models (LLMs) advances, it addresses numerous practical challenges while raising many new issues. In particular, the widespread generation of text by LLMs on the Internet may increase the spread of rumors and raise concerns about text copyright (Megías et al. 2021; Tang et al. 2023). Consequently, the identification and classification of machine-generated text have become critically significant. Watermarking techniques for LLMs can help to tackle these problems, leading to their rising importance in ongoing conversations and attracting increasing interest globally.
Text watermarking involves embedding distinctive, imperceptible identifiers (watermarks) into written content. Nowadays, most methods are logit-based (Kirchenbauer et al. 2023a; Liu et al. 2023a; Zhao et al. 2023; Kuditipudi et al. 2023; Hu et al. 2023): they manipulate the output logits of LLMs during the generation process using distinct but consistently logit-based strategies to embed watermarks successfully. Utilizing the power of LLMs ensures that such adjustings for probabilistic distribution are seamlessly integrated into the generated content without compromising the overall quality or coherence of the text (Lin et al. 2024). These watermark methods are sophisticatedly designed to be robust yet discreet, ensuring content integrity and ownership preservation without compromising readability or meaning.
@@ Introduction
However, as watermarking techniques proliferate, watermark collision becomes inevitable with the increasing application of watermarks. The term watermark collision can be defined as instances where the text contains multiple watermarks simultaneously (Fig. 1). This is particularly inevitable during tasks that may require the collaboration of multiple LLMs, such as paraphrasing and translation. While many methods (Liu et al. 2023a; Zhao et al. 2023; Kuditipudi et al. 2023) claim resilience against paraphrase attacks, none have been specifically tested for watermark collisions. Hence, we examine the underlying mathematical principles, employ watermark collision as an attack strategy for all watermarking techniques that utilize logits, and evaluate its effectiveness in conjunction with multiple traditional attack methods enhanced through the incorporation of watermark collision.
@@
Modern text watermarking techniques can be classified into two categories: modification-based and generation-based watermarks (Liu et al. 2024). Modification-based watermarking, also known as watermarks for existing text, consists of altering an existing text to create a watermarked text. Most of the modification-based techniques can be classified as either lexical (Topkara, Topkara, and Atallah 2006b; Yang et al. 2022; Munyer and Zhong 2023; Sato et al. 2023) or syntactic methods (Atallah et al. 2001; Topkara, Topkara, and Atallah 2006a; Meral et al. 2009), based on rules, classical machine learning or deep neural models.
@@
While modification-based techniques (Abdelnabi and Fritz 2020; Yang et al. 2022) modify the text and preserve its semantics, generation-based methods apply watermarks into the text generation process to achieve better results, enabling smoother integration with LLMs. Watermark injection can be carried out during either the training phase or the inference phase.
During the training of LLMs, watermarks are inserted into the training data to intentionally alter the results of LLMs for certain inputs (Liu et al. 2023b; Sun et al. 2022). The main objective of training time watermarking is to protect dataset copyrights from unauthorized usage (Tang et al. 2023; Sun et al. 2023). Despite being able to embed watermarks in LLMs, training-time watermarking has significant limitations, including limited payload capacity, restricted trigger conditions, and significant training overhead.
For inference-time watermarking, Kirchenbauer et al. (2023a) proposed a logit-based greenlist mechanism based on prior token hashes. Liu et al. (2023a) introduced a watermark model to generate semantic-preserving logits during text generation. Zhao et al. (2023) simplified the Kirchenbauer et al.’s scheme by using a fixed Green-Red split and achieved greater robustness. Christ, Gunn, and Zamir (2023); Kuditipudi et al. (2023); Fu et al. (2024) aim to design watermark techniques that are more robust and secure. Yoo, Ahn, and Kwak (2023); Boroujeny et al. (2024) introduce watermarking techniques with increased payload capacity for arbitrary binary data. Zhu et al. (2024) enhances the efficiency and quality of watermarking by embedding dual secret patterns.
Even though these methods have been designed to be more robust against attacks such as paraphrase attacks (Kirchenbauer et al. 2023b), back-translation attacks (He et al. 2024) and mask-and-fill attacks (Lyu, Huang, and Kong 2023), these attacks often use unwatermarked LLMs, e.g. DIPPER (Krishna et al. 2023) and GPT-3.5 (Brown et al. 2020). Prior research in the pre-LLM era has mentioned potential risks associated with multiple watermarks (Tanha et al. 2012). Nevertheless, the effects of one watermarking technique on another in the context of LLM watermarking remain unclear, which is the motivation for this study.
Method
@@
Principle of Watermark Collision
The detection process for logit-based watermarking methods relies on the null hypothesis testing. A well-known example is the null hypothesis of KGW (Kirchenbauer et al. 2023a):
@@
Let $\mathbf{P}(T)$ represent the probability distribution of a certain text $T$. The detection process verifies if $\mathbf{P}(T)$ follows the null hypothesis $H_{0}$, which can be formulated as the follows ($\bm{\delta}(\cdot)$ is the dirichlet function):
@@
Since the detection process relies on identifying patterns consistent with $\mathbf{P}_{w^{(i)}}$, the introduction of $\mathbf{P}_{\text{entangled}}$ causes the final text distribution to no longer conform to any of these patterns. The detection algorithm may not be able to detect the watermark in the text due to a shift in word distribution from $P_{w^{(i)}}$, leading to watermark collisions.
Pipeline Design
@@
Watermarker
Watermarker $W$ generates watermarked texts $T_{W}$ by using a language model (LM) to create content based on a specific corpus as context. As illustrated in Fig. 2, we first produce the watermarked text data $T_{W}$ with Watermarker $W$. Additionally, we generate unwatermarked text $T_{W^{\prime}}$ using the same context and prompt as $T_{W}$ for further comparisons. Experiments Section and Appendix provide details regarding the watermarker setup.
Colliders
@@
It is important to note that the KGW-based method is essentially ineffective against back-translation attacks due to the incapability of capturing the contextual semantics. Thus, they have been excluded from this pipeline. We choose French as the pivot language in the back-translation. Details are in Experiments Section.
@@
Detectors
As demonstrated in Fig.2, there are four detectors, each tailored to identify a specific type of watermark. Detector $D_{P}$ targets watermarks in paraphrasers, $D_{R}$ focuses on those in translators, and $D_{M}$ is for watermarks in the MnF process. Detector $D_{W}$ aims to identify the original watermark embedded by the watermarker. By comparing the results from these detectors, we can assess the effectiveness of the attacks with or without additional watermarks.
@@
Experiments
Experiment Setup
@@
We utilize the C4 dataset (Raffel et al. 2020) as the context for text generation with a maximum of 128 tokens. For watermarker and paraphraser, we employ LLaMA-2-13B (Touvron et al. 2023), Qwen2-7B (Bai et al. 2023) and OPT-1.3B (Zhang et al. 2022). For back-translator, we use LLaMA-2-13B and Qwen2-7B. For mask-and-filler, we exclusively utilize RoBERTaLARGE (Liu et al. 2019) as the masked language model. We only present results of LLaMA-2-13B here and refer to the Appendix for more details of the experiments.
@@
Evaluation Metrics
Given the rapid development of the watermarking field, there is currently no consensus on metrics, with different methods employing varied evaluation criteria (Tu et al. 2023). Moreover, some approaches utilize a detection threshold to categorize text as watermarked, while others differ. Implementing all possible metrics for each method is impractical and biased.
Following the approach of Zhao et al. (2023), we opt for a fair evaluation by avoiding the influence of threshold settings. Our final detection metrics include false positive rates (FPR) and true positive rates (TPR). We specifically set FPR values at 1%, 5%, and 10%, adjusting the detector’s thresholds accordingly. This ensures a consistent and unbiased assessment of watermarking methods. We only present results when $\text{FPR}=1\%$ here and refer to Appendix for comprehensive results.
@@
Table 4: PPL of the back-translated $T_{P}$ with dual WMs.
Experimental Results & Analysis
@@
As shown in Tab. 3 and 4, we present perplexity before and after attacks, both with and without collisions. We use LLaMA-2-13B as the backbone for perplexity calculation, which is the same model for the initial generated text. As evidence, the text quality remains largely stable post-attack, and most collisions did not result in significant declines in text quality, indicating the potential value of collision as a methodology of attack.
@@
However, some anomalies can be observed in back-translation (Table 2), where the TPR increases after a collision. These anomalies occur because certain watermarks may have similar distributions in specific contexts, rendering them ineffective and making attacks on them pointless: When two watermark methods exhibit similar distributions, the likelihood of collision decreases, leading to an increase in the True Positive Rate (TPR). For example, if two KGW-based watermark methods utilize the a similar but not the same red-green list, paraphrasing words into other similar words may accidentally reinforce the watermark within the text, thereby enhancing the TPR. But this is a degraded performance for a certain watermark method: It is a challenge to identify the originating method of the watermark. In practical scenarios, if the origin of the watermark cannot be determined, it essentially means we cannot ascertain which entity embedded the watermark. Consequently, such cases are considered failures of specific watermark methods and do not concern us.
@@
We note that attacks on MnF are less effective because the unmasked words retain most of the context and keep the watermark unaltered. Therefore, the TPR in these cases remains nearly the same. Nonetheless, it is observed that when considering the z-score of $H_{0}$ as stated in (Kirchenbauer et al. 2023a), the influence of watermark collision persists.
In Table 5, we present the z-scores for each corresponding method. Degradation of the z-score is observed after collision attacks compared to those without, indicating the effectiveness of collisions. Besides, the mask rate has a greater impact on the detection process. If the mask rate is low, MnF is less likely to have a significant effect, thus explaining why these attacks have minimal impact on MnF. Furthermore, the SIR method is excluded from MnF experiments, as it prioritizes semantic factors and is less susceptible to MnF attacks.
@@
Multi-round Collision
@@
Possible Application
Interference with downstream tasks caused by watermark collisions.
Since LLM watermarks may be used in a variety of downstream applications, particularly in QA scenarios, collisions between watermarks may cause greater damage to downstream watermarks. Texts with watermarks are more likely to be used as user input or dialogue context. As a result, LLM service providers may have difficulty watermarking their output. This may occur more frequently and harm the development of LLM watermarking techniques.
@@
Previous works (Kirchenbauer et al. 2023b, a; Kuditipudi et al. 2023) have introduced several attacks, such as copy-paste attacks and paraphrase attacks, but most have shown their robustness and security against at least some of these attacks. Our study, however, provides a feasible method of constructing effective attacks using watermark collisions. For example, text with a KGWwk watermark can be detected with a 44.8% TPR after a paraphrase attack without a watermark. However, if the paraphrase attack is conducted with another KGWwk, the detection TPR drops to 41.9%. If the attack is done with a KGWsg, the detection TPR further decreases to 28.5%, as presented in Table 1(a) and 1(b). The use of paraphrase, back-translation, and mask-and-fill colliders with strong watermarks could easily wipe out existing watermarks, resulting in greater vulnerability to watermarking.
@@
In Table 1 and 2, we demonstrate that weak watermarks can still be easily applied to unwatermarked text. It is, however, much more difficult to apply a weak watermark to text that has already been watermarked (Tab. 1 & 2). For example, KGW can be applied to plaintext with a success rate of 99.90%, but can be applied to a SIR-watermarked text with a probability not exceeding 25.10%, as presented in Table 1(a) and 1(c). This provides a simple probabilistic method of detecting watermarks without the need to know their details. When it is difficult to add a weak watermark to a paraphrased sentence, it is more likely to have an existing watermark in the original sentence.
Discussion
@@
  • Black-box Detection: As we discussed in the Experiments Section, the watermark collisions could be used to perform black-box detection of any existing watermarks in any text. Users would experience distrust when they become aware of the presence of watermarks. Hackers may attempt to bypass the watermarks.
Conclusion
In this study, we examine how overlapping watermarks in the same text can decrease the accuracy of both upstream and downstream watermark detection. We propose the use of watermark collisions as an attacking philosophy and therefore emphasize that watermark collisions may compromise the validity and security of all logit-based watermarks. We conduct experiments by integrating various watermarkers and colliders, assessing the text quality before and after collisions, as well as with and without collisions. Our results demonstrate that collision can enhance common attacks and text quality remains consistent when compared to attacks without collisions. While we do not aim to solve the issue of watermark collisions, we hope our work will increase awareness of potential threats to LLM watermarking.

### v3

@@ Abstract
Lost in Overlap: Exploring Logit-based Watermark Collision in LLMs
Yiyang Luo* Nanyang Technological University lawrenceluoyy@outlook.com                        Ke Lin* Tsinghua University leonard.keilin@gmail.com                        Chao Gu* University of Science and Technology of China guch8017@mail.ustc.edu.cn
Jiahui Hou University of Science and Technology of China jhhou@ustc.edu.com                        Lijie Wen Tsinghua University wenlj@tsinghua.edu.cn                        Ping Luo Tsinghua University luop@tsinghua.edu.cn
footnotetext: *Equal Contribution.footnotetext: Code and data are available at https://github.com/AInnovateLab/watermark-collision.
1 Introduction
As the quality of text produced by large language models (LLMs) advances, it addresses numerous practical challenges while raising many new issues. In particular, the widespread generation of text by LLMs on the Internet may increase the spread of rumors and raise concerns about text copyright Megías et al. (2021); Tang et al. (2023). Consequently, the identification and classification of machine-generated text have become critically significant. Watermarking techniques for LLMs can help to tackle these problems, leading to their rising importance in ongoing conversations and attracting increasing interest globally.
Text watermarking involves embedding distinctive, imperceptible identifiers (watermarks) into written content. Nowadays, most methods are logit-based Kirchenbauer et al. (2023a); Liu et al. (2023a); Zhao et al. (2023); Kuditipudi et al. (2023); Hu et al. (2023): they manipulate the output logits of LLMs during the generation process using distinct but consistently logit-based strategies to embed watermarks successfully. Utilizing the power of LLMs ensures that such adjustings for probabilistic distribution are seamlessly integrated into the generated content without compromising the overall quality or coherence of the text Lin et al. (2024). These watermark methods are sophisticatedly designed to be robust yet discreet, ensuring content integrity and ownership preservation without compromising readability or meaning.
@@ 1 Introduction
However, as watermarking techniques proliferate, watermark collision becomes inevitable with the increasing application of watermarks. The term watermark collision can be defined as instances where the text contains multiple watermarks simultaneously (Fig. 1). This is particularly inevitable during tasks that may require the collaboration of multiple LLMs, such as paraphrasing and translation. While many methods Liu et al. (2023a); Zhao et al. (2023); Kuditipudi et al. (2023) claim resilience against paraphrase attacks, none have been specifically tested for watermark collisions. Hence, we examine the underlying mathematical principles, employ watermark collision as an attack strategy for all watermarking techniques that utilize logits, and evaluate its effectiveness in conjunction with multiple traditional attack methods enhanced through the incorporation of watermark collision.
@@ 2 Related Work
Modern text watermarking techniques can be classified into two categories: modification-based and generation-based watermarks Liu et al. (2024). Modification-based watermarking, also known as watermarks for existing text, consists of altering an existing text to create a watermarked text. Most of the modification-based techniques can be classified as either lexical Topkara et al. (2006b); Yang et al. (2022); Munyer and Zhong (2023); Sato et al. (2023) or syntactic methods Atallah et al. (2001); Topkara et al. (2006a); Meral et al. (2009), based on rules, classical machine learning or deep neural models.
@@
While modification-based techniques Abdelnabi and Fritz (2020); Yang et al. (2022) modify the text and preserve its semantics, generation-based methods apply watermarks into the text generation process to achieve better results, enabling smoother integration with LLMs. Watermark injection can be carried out during either the training phase or the inference phase.
During the training of LLMs, watermarks are inserted into the training data to intentionally alter the results of LLMs for certain inputs Liu et al. (2023b); Sun et al. (2022). The main objective of training time watermarking is to protect dataset copyrights from unauthorized usage Tang et al. (2023); Sun et al. (2023). Despite being able to embed watermarks in LLMs, training-time watermarking has significant limitations, including limited payload capacity, restricted trigger conditions, and significant training overhead.
For inference-time watermarking, Kirchenbauer et al. (2023a) proposed a logit-based greenlist mechanism based on prior token hashes. Liu et al. (2023a) introduced a watermark model to generate semantic-preserving logits during text generation. Zhao et al. (2023) simplified the Kirchenbauer et al.’s scheme by using a fixed Green-Red split and achieved greater robustness. Christ et al. (2023); Kuditipudi et al. (2023); Fu et al. (2024) aim to design watermark techniques that are more robust and secure. Yoo et al. (2023); Boroujeny et al. (2024) introduce watermarking techniques with increased payload capacity for arbitrary binary data. Zhu et al. (2024) enhances the efficiency and quality of watermarking by embedding dual secret patterns.
Even though these methods have been designed to be more robust against attacks such as paraphrase attacks Kirchenbauer et al. (2023b), back-translation attacks He et al. (2024) and mask-and-fill attacks Lyu et al. (2023), these attacks often use unwatermarked LLMs, e.g. DIPPER Krishna et al. (2023) and GPT-3.5 Brown et al. (2020). Prior research in the pre-LLM era has mentioned potential risks associated with multiple watermarks Tanha et al. (2012). Nevertheless, the effects of one watermarking technique on another in the context of LLM watermarking remain unclear, which is the motivation for this study.
3 Method
@@
3.1 Principle of Watermark Collision
The detection process for logit-based watermarking methods relies on the null hypothesis testing. A well-known example is the null hypothesis of KGW Kirchenbauer et al. (2023a):
@@ 3.1 Principle of Watermark Collision
Let $\mathbf{P}(T)$ represent the probability distribution of a certain text $T$. The detection process verifies if $\mathbf{P}(T)$ follows the null hypothesis $H_{0}$, which can be formulated as the follows, where $\bm{\delta}(\cdot)$ is the dirichlet function:
@@
Since the detection process relies on identifying patterns consistent with $\mathbf{P}_{w^{(i)}}$, the introduction of $\mathbf{P}_{\text{entangled}}$ causes the final text distribution to no longer conform to any of these patterns. The detection algorithm may not be able to detect the watermark in the text due to a shift in word distribution from $P_{w^{(i)}}$, leading to watermark collisions, whereas standard attacks are random, unstable, and less effective.
3.2 Pipeline Design
@@
3.2.1 Watermarker
Watermarker $W$ generates watermarked texts $T_{W}$ by using a language model (LM) to create content based on a specific corpus as context. As illustrated in Fig. 2, we first produce the watermarked text data $T_{W}$ with Watermarker $W$. Additionally, we generate unwatermarked text $T_{W^{\prime}}$ using the same context and prompt as $T_{W}$ for further comparisons. Section 4 and Appendix A provide details regarding the watermarker setup.
3.2.2 Colliders
@@
It is important to note that the KGW-based method is essentially ineffective against back-translation attacks due to the inability to capture the contextual semantics. Thus, they have been excluded from this pipeline. We choose French as the pivot language in the back-translation. Details are in Section 4.
@@
3.2.3 Detectors
As demonstrated in Fig.2, four detectors are tailored to identify a specific type of watermark. Detector $D_{P}$ targets watermarks in paraphrasers, $D_{R}$ focuses on those in translators, and $D_{M}$ is for watermarks in the MnF process. Detector $D_{W}$ aims to identify the original watermark embedded by the watermarker. By comparing the results from these detectors, we can assess the effectiveness of the attacks with or without additional watermarks.
@@ 3.2.3 Detectors
4 Experiments
4.1 Experiment Setup
@@
We utilize the C4 dataset Raffel et al. (2020) as the context for text generation with a maximum of 128 tokens. For watermarker and paraphraser, we employ LLaMA-2-13B Touvron et al. (2023), Qwen2-7B Bai et al. (2023) and OPT-1.3B Zhang et al. (2022). For back-translator, we use LLaMA-2-13B and Qwen2-7B. For mask-and-filler, we exclusively utilize RoBERTaLARGE Liu et al. (2019) as the masked language model. We only present the results of LLaMA-2-13B here and refer to the Appendix B for more details.
@@
4.2 Evaluation Metrics
Given the rapid development of the watermarking field, there is currently no consensus on metrics, with different methods employing varied evaluation criteria Tu et al. (2023). Moreover, some approaches utilize a detection threshold to categorize text as watermarked, while others differ. Implementing all possible metrics for each method is impractical and biased.
Following the approach of Zhao et al. (2023), we opt for a fair evaluation by avoiding the influence of threshold settings. Our final detection metrics include false positive rates (FPR) and true positive rates (TPR). We specifically set FPR values at 1%, 5%, and 10%, adjusting the detector’s thresholds accordingly. This ensures a consistent and unbiased assessment of watermarking methods. We only present results when $\text{FPR}=1\%$ here and refer to Appendix B for comprehensive results.
@@ 4.2 Evaluation Metrics
Table 4: PPL of the back-translated $T_{R}$ with dual WMs.
4.3 Experimental Results & Analysis
@@
As shown in Tab. 3 and 4, we present perplexity before and after attacks, both with and without collisions. We use LLaMA-2-13B as the backbone for perplexity calculation, which is the same model for the initial generated text. As evidence, text quality remains largely stable post-attack, and most collisions did not result in significant declines in text quality, indicating the potential value of collision as an attack methodology. Detailed semantic analysis examples of the text quality can be found in Appendix B.
@@
However, some anomalies can be observed in back-translation (Table 2), where the TPR increases after a collision. These anomalies occur because certain watermarks may have similar distributions in specific contexts, rendering them ineffective and making attacks on them pointless: When two watermark methods exhibit similar distributions, the likelihood of collision decreases, leading to an increase in the True Positive Rate (TPR). However, in fact, this is a degraded performance for a certain watermark method: It is a challenge to identify the origin of the watermark. In practical scenarios, if the origin of the watermark cannot be determined, it essentially means we cannot ascertain which entity embeds the watermark and makes the watermark meaningless. For example, if two KGW-based watermark methods utilize the a similar but not the same red-green list, paraphrasing words into other similar words may accidentally reinforce the watermark within the text, thereby enhancing the TPR for both watermarks and lead to confusion on which entity apply watermark on the text. Consequently, such cases are considered failures of watermark methods and do not concern us.
@@
We note that attacks on MnF are less effective because the unmasked words retain most of the context and keep the watermark unaltered. Therefore, the TPR in these cases remains nearly the same.
Nonetheless, it is observed that when considering the z-score of $H_{0}$ as stated in Kirchenbauer et al. (2023a), the influence of watermark collision persists. In Table 5, we present the z-scores for each corresponding method. Degradation of the z-score is observed after collision attacks compared to those without, indicating the effectiveness of collisions. Besides, the mask rate has a greater impact on the detection process. If the mask rate is low, MnF is less likely to have a significant effect, thus explaining why these attacks have minimal impact on MnF. Furthermore, the SIR method is excluded from MnF experiments, as it prioritizes semantic factors and is less susceptible to MnF attacks.
@@
4.4 Multi-round Collision
@@
5 Possible Application
@@
Previous works Kirchenbauer et al. (2023b, a); Kuditipudi et al. (2023) have introduced several attacks, such as copy-paste attacks and paraphrase attacks, but most have shown their robustness and security against at least some of these attacks. Our study, however, provides a feasible method of constructing effective attacks using watermark collisions. For example, text with a KGWwk watermark can be detected with a 44.8% TPR after a paraphrase attack without a watermark. However, if the paraphrase attack is conducted with another KGWwk, the detection TPR drops to 41.9%. If the attack is done with a KGWsg, the detection TPR further decreases to 28.5%, as presented in Table 1(a) and 1(b). The use of colliders with strong watermarks could easily erase existing watermarks, resulting in greater vulnerability to watermarking.
@@
In Table 1 and 2, we demonstrate that weak watermarks can still be easily applied to unwatermarked text. It is, however, much more difficult to apply a weak watermark to text that has already been watermarked (Tab. 1 & 2). For example, KGW can be applied to plaintext with a success rate of 99.90%, but can be applied to a SIR-watermarked text with a probability not exceeding 25.10%, as presented in Table 1(a) and 1(c). This provides a simple probabilistic method of detecting watermarks without the need to know their details. When adding a weak watermark to a sentence is difficult, it is more likely to have an existing one in the original sentence.
6 Discussion
@@
  • Black-box Detection: As we discussed in the Experiments Section, the watermark collisions could perform black-box detection of any existing watermarks in any text. Users would experience distrust when they become aware of the presence of watermarks. Hackers may attempt to bypass the watermarks.
7 Conclusion
In this study, we examine how overlapping watermarks in the same text can decrease the accuracy of both upstream and downstream watermark detection. We propose the use of watermark collisions as an attacking philosophy and therefore emphasize that watermark collisions may compromise the validity and security of all logit-based watermarks. We conduct experiments by integrating various watermarkers and colliders, assessing the text quality before and after collisions, as well as with and without collisions, to demonstrate that collision can enhance common attacks and text quality remains consistent when compared to attacks without collisions. We hope our work will increase awareness of potential threats to LLM watermarking.
Limitations
Our approach indeed demonstrates the potential collision between existing watermark techniques. Nevertheless, we conduct our experiments only on paraphrasing, back-translation, and mask-and-fill tasks to simulate watermark collisions. Broader but less related tasks, such as question answering, have not yet been tested.
Furthermore, the range of models chosen is limited. The models chosen include LLaMA-2-13B, Qwen2-7B, OPT-1.3B, and RoBERTaLARGE as the colliders, while certain other models are tailored for particular tasks. Experiments on more open-sourced models could further enhance the conclusion of our paper.